### v1

# Autoencoded Image Compression for Secure and Fast Transmission

Abstract
With an exponential growth in the use of digital image data, the need for efficient transmission methods has become imperative. Traditional image compression techniques often sacrifice image fidelity for reduced file sizes, presenting a challenge in maintaining both quality and efficiency. They also tend to compromise on security, leaving images vulnerable to threats such as man-in-the-middle attacks. This paper proposes an autoencoder architecture for image compression so as to not only help in dimensionality reduction but also inherently encrypt the images. The paper also introduces the use of a composite loss function that combines reconstruction loss and residual loss for improved performance. The autoencoder architecture is designed to achieve optimal dimensionality reduction and regeneration accuracy while safeguarding the compressed data during transmission or storage. Images regenerated by the autoencoder are evaluated against three key metrics: reconstruction quality, compression ratio, and one-way delay during image transfer. The experiments reveal that the proposed architecture achieves an SSIM of 97.5% over the regenerated images and an average latency reduction of 87.5%, indicating its effectiveness as a secure and efficient solution for compressed image transfer.
Index Terms:
Autoencoder, image compression, latent space, secure transmission, Structural Similarity Index Measure (SSIM), latency reduction
I Introduction
Convolutional Autoencoder is a type of neural network architecture specifically designed to learn latent representations of data offering a powerful mechanism for unsupervised learning and data compression. In this paper, we introduce a novel method of utilising autoencoders for compressed image transfer. We also explore the usage of a residual component added to the loss function to improve the quality of the reconstructed image. A convolutional autoencoder mainly comprises two key components:
Encoder: It consists of a convolutional neural network that takes an image as input and progressively extracts its key features through convolution and pooling layers. These features are then encoded into a lower-dimensional latent space representation, capturing the essence of the image in a significantly smaller size.
Decoder: It consists of a mirrored CNN architecture that takes the latent space representation generated by the encoder and upsamples it through deconvolutional layers. The decoder progressively reconstructs the image, aiming to recreate the original image as accurately as possible.
Figure 1: Autoencoder architecture
The latent space representation generated by the encoder holds the key to inherent encryption. Unlike traditional compression techniques that operate directly on the image data, an autoencoder [1] compresses the image into a latent space representation that is not directly interpretable. The compressed representation is a product of the complex non-linear transformations learned by the encoder’s convolutional layers. Decoding this representation requires the specific architecture and weights of the trained decoder, making it inherently encrypted for anyone without access to the decoder network.
This paper presents an approach to secure image transmission [2] utilising a convolutional autoencoder trained on a designated image dataset, allowing it to learn the optimal representation of the images. Following training, the components of the autoencoder are separated. The encoder resides at the sender’s host, while the decoder resides at the receiver’s host. Images are compressed using the sender-side encoder, generating a latent space representation that is significantly smaller than the original image. The compressed data is then transmitted across the network using a client-server architecture. Using the above architecture, two key advantages are noticeable: The transmission times is significantly less compared to traditional methods, and the image is seemingly encrypted. The latent space representation itself acts as inherent encryption. Without the decoder present on the receiver’s side, the compressed data is unintelligible, effectively preventing unauthorised access or manipulation during transmission. Hence, the distributed architecture offers a secure and efficient solution for image transmission.
The paper is structured as follows: Section II provides a literature review on the application of autoencoders for image compression as proposed in existing research. Section III outlines the datasets used for training and evaluation while section IV delves into the architecture of the autoencoder proposed in this study. Section V elaborates on the training process of the autoencoder and introduces the novel loss function proposed. Section VI presents the obtained results and compares them with JPEG compression. Lastly, Section VII offers concluding remarks and discusses potential future directions.
II Literature review
Autoencoders have emerged as a promising tool for various tasks, including dimensionality reduction [3] and image compression [4]. Leveraging deep learning techniques, researchers have explored innovative approaches to enhance image retrieval and compression algorithms. Theis et al. [5] introduced a novel approach combining deep learning with autoencoders to ensure better image retrieval, which was evaluated using basic distance functions. They concluded that increasing the number of dimensions in fused features led to better retrieval performance. Following this principle, increasing the number of filters to double after each layer allows for the capture of complex and abstract features.
Cheng et al. [6] proposed a deep convolutional autoencoder-based image compression method that significantly reduced the bit rate compared to JPEG, demonstrating superior compression. Their approach integrated Principal Component Analysis (PCA) to rotate feature maps, ensuring an energy-compact representation that enhanced entropy coding. This method handled image sizes up to $256\times 256$, leading to notable improvements in the quality of reconstructed images.
Additionally, Petscharnig et al. [7] proposed lossy image compression using compressive autoencoders, achieving performance that often matched or exceeded JPEG2000 in terms of quality, as verified by PSNR, SSIM, and MS-SSIM. Their approach showed the potential to adapt quickly to changing tasks due to its neural network architecture, though it required complex training. They concluded that compressive autoencoders produced good performance, often outperforming traditional methods like JPEG2000 in terms of SSIM and MS-SSIM while maintaining PSNR.
By combining these insights and advancements, the overall concept of using autoencoders for image compression and secure image transfer can be significantly enhanced.
III Dataset
III-A Training Dataset
The Stanford Dogs [8] and Animals 10 [9] datasets were used for training and evaluating the proposed convolutional autoencoder model. The Stanford Dogs dataset comprises 20,580 color images spanning 120 different dog breeds. These images were initially gathered from the ImageNet [10] dataset. The Animals 10 dataset comprises 26,200 images belonging to 10 categories: dog, cow, horse, elephant, butterfly, cat, chicken, sheep, spider, and squirrel. These were obtained from Google Images.
The Stanford Dogs dataset provides annotations, including class labels specifying the dog breed for each image and bounding box coordinates locating the dog within each image. However, for this work on image compression, only the image data itself was used, while bounding box annotations were not utilized.
The images were preprocessed for ensuring compatibility with the convolutional autoencoder model. Each image was resized to a fixed resolution of $256\times 256$ pixels in RGB format, while preserving the original aspect ratio using bicubic interpolation. This resizing was necessary to meet the neural network’s requirement of consistent input sizes during both training and inference. The original images in the datasets had varying resolutions, with most falling in the range of 200 to 500 pixels along the larger dimension.
After resizing, the images were encoded as tensors with a shape of (256, 256, 3), representing height, width, and color channels (RGB), respectively. The following data augmentations were applied to the training dataset:
  • Random Horizontal Flip
  • Random Vertical Flip
  • Random Rotation through 15 degrees
III-B Evaluation Dataset
For evaluation, we used the Natural Images dataset [11], which consists of images from various classes. Eight images were selected, one from each class, for testing.
IV Architecture
The proposed architecture consists of a convolutional autoencoder designed for progressive image compression. The encoder encodes the input image into a latent space representation, while the decoder aims to reconstruct the original image from this compressed form. The key novelty of our approach lies in the loss calculation, which combines residual loss and reconstruction loss. The residual loss, inspired by image super-resolution models [12] [13], captures high-frequency details. Minimizing the residual loss enables the model to quickly learn efficient representations. Meanwhile, the reconstruction loss encourages the autoencoder to develop a compact representation of the input image, facilitating effective compression. By jointly optimizing these losses, our model achieves an optimal trade-off between image quality and compression ratio. Existing methods typically rely on entropy coding and quantization for compression. However, our approach achieves comparable compression results using solely an autoencoder architecture. The autoencoder learns to compress and reconstruct images in an end-to-end fashion, eliminating the need for traditional compression techniques.
Figure 2: Model setup during training
Figure 3: Model setup during testing
IV-A Encoder
The encoder comprises of four convolutional layers with ReLU activations, each succeeded by a $2\times 2$ max-pooling operation for downsampling. The number of filters doubles after each convolutional layer, beginning with 32 in the first layer and reaching 256 in the fourth layer. This progressive increase in filters enables the encoder to capture increasingly complex and abstract features with greater depth. Lastly, a bottleneck convolutional layer having 64 filters is employed to transform the encoded representation to a lower-dimensional space.
IV-B Decoder
The decoder architecture is symmetrical to the encoder. It has four transpose convolutional (deconvolution) layers with ReLU activations to progressively upsample and reconstruct the original image. The number of filters in the decoder layers decreases by a factor of 2 after each layer, starting from 256 and ending with 32 filters. The final layer is a transpose convolutional layer with 3 filters and a sigmoid activation to produce the reconstructed image in the RGB colour space.
The progressive compression aspect of the model is achieved by allowing the decoder to output not only the reconstructed image but also a residual image during training, as shown in Figure 2. This residual image represents the difference between the original input image and the reconstructed image, capturing the information lost during the compression and reconstruction process. Throughout training, the model minimizes the information carried by the residual image. During inference, the model wrapper responsible for outputting the residual image is removed.
V Training
V-A Model and Setup
The training process employed a progressive compression method, integrating reconstruction and residual components to optimize the convolutional autoencoder for image compression. The encoder and decoder components were encapsulated within a model wrapper, allowing the model to output both the reconstructed and residual images during training. The residual image, representing the difference between the input and reconstructed images, facilitated higher reconstruction quality and reduced training times.
Training was conducted on two NVIDIA T4 GPUs, leveraging PyTorch’s DataParallel module for parallel computation across both GPUs. This parallelization enabled efficient training with larger batch sizes.
[…]
$d$ $=\text{Decoder}(e)$
$r$ $=x-d$
$(d,r)$ $=\text{Model}(x)$
V-B Hyperparameters
Our experiment revealed that a batch size of 8 yielded the best results, as detailed in the results section. We used an 80/20 train/validation split. The Adam optimizer with an initial learning rate of 0.001 was utilized. To ensure convergence and prevent overfitting, a ’reduce learning rate on plateau’ scheduler was used. This scheduler reduced the learning rate by a factor of 0.1 when the validation loss remained stagnant for 10 successive epochs. Training was conducted for a total of 100 epochs. Additionally, the Early stopping mechanism halted training if the validation loss failed to improve over 20 epochs, thereby preventing overfitting.
V-C Loss Calculation
The objective loss function guiding the training process was the sum of the reconstruction loss and residual loss, each calculated by the mean squared error (MSE). The reconstruction loss measured the difference between the input image and the image reconstructed by the decoder. The residual loss quantified the difference between the residual image and a target tensor of zeros. By minimizing both the reconstruction error and the residual image, the final loss used for training enabled the model to learn an efficient compression strategy.
Figure 4: Loss vs Epochs during training
[…]
$\mathcal{L}_{i}$ $=\text{Criterion}(r,\mathbf{0})$
$\mathcal{L}$ $=\mathcal{L}_{r}+\mathcal{L}_{i}$
Key:
[…]
$\mathcal{L}$ $:\text{Total loss}(\mathcal{L}_{r}+\mathcal{L}_{i})$
VI Results
VI-A Effect of the Residual Component
The architecture was evaluated by comparing the performance of two models: one incorporating residual loss and the other without it. Both models shared the same architecture; however, during the forward pass, one model additionally produced a residual image.
To expedite the training process, a batch size of 256 and Distributed Data Parallel (DDP) were utilized. The models were trained for 20 epochs using the training setup detailed in the previous section.
TABLE I: Comparison between models with and without progressive compression
Model PSNR SSIM MSE
Without PgIC 29.5916 0.8869 0.0012
With PgIC (proposed) 31.3478 0.8969 0.0008
After training, the models were evaluated using 8 test images from the natural images dataset, each representing a different class. The models were used to encode and decode these images. Results indicated that the model optimized with residual loss achieved higher Peak Signal to Noise Ratio (PSNR) and Structural Similarity Index (SSIM)[14], and lower Mean Square Error (MSE). These metrics collectively demonstrate that the model with residual loss provided better overall reconstruction quality.
(a) Results with PgIC
(b) Results without PgIC
Figure 5: Comparison of results with and without PgIC
VI-B Performance Evaluation Across Varied Batch Sizes using PgIC
TABLE II: Performance evaluation for different batch sizes
Batch Size PSNR SSIM MSE
8 37.0432 0.9757 0.0002
16 36.9585 0.9748 0.0002
32 37.0554 0.9754 0.0002
64 36.2745 0.9708 0.0003
128 34.8692 0.9651 0.0004
256 32.4127 0.9295 0.0006
Selecting the optimal batch size is pivotal in training deep learning models. Thus, we conducted experiments across a range of batch sizes, from 8 to 256. The evaluation for the optimal batch size was performed using the PgIC. Each model underwent training for 100 epochs using the training dataset. Subsequently, they were assessed in a testing environment using a set of 64 images sourced from the natural images dataset, with 8 images sampled from each of the 8 classes. It was observed that the optimal performance was achieved with a batch size of 8, yielding an SSIM of 97.57%.
VI-C Compression Ratio and Transfer Time
Figure 6: One-way transmission latency of compressed and uncompressed images
Each input image was resized to a dimension of (256,256,3) to ensure compatibility with the encoder architecture. The compressed latent space had a dimension of (16,16,64). Thus the architecture was able to achieve a compression ratio 12:1 with respect to the resized images.
To evaluate the effectiveness of the compression in a networked environment, we established a setup with two hosts connected via TCP sockets. Under identical network conditions, we transferred both the compressed and uncompressed images from one host to the other and recorded the one-way latency. The results of these transfers are illustrated in Figure 6.The evaluation revealed an average reduction in latency from uncompressed to compressed images to be 87.5%.
VI-D Comparison with exisiting methods
Existing image compression methods typically achieve a compression ratio of about 2:1 while maintaining a high Structural Similarity Index (SSIM) of around 98% to 99%, ensuring minimal loss in image quality. In contrast, our proposed autoencoder not only achieves a comparable SSIM of 97.5%, but also significantly outperforms traditional methods with a compression ratio of 12:1. Additionally, the autoencoder incorporates encryption, providing enhanced security for the compressed images. We compared our autoencoder to the widely used JPEG algorithm across various quality settings. We analyzed the trade-off between compression ratio and image quality metrics (PSNR and SSIM) for JPEG quality levels (10-100). The results show that our autoencoder achieves comparable PSNR and SSIM values to JPEG at a quality setting of 75, but with a 6x higher compression ratio. The detailed comparison is illustrated in Figure 7. This superior compression efficiency, combined with the inherent encryption capabilities, makes our autoencoder ideal for applications requiring both secure data transmission and efficient storage, all without sacrificing visual quality.
(a) SSIM vs compression ratio
(b) PSNR vs compression ratio
Figure 7: SSIM and PSNR vs compression ratio values for various image compression techniques
VII Conclusion and Future Work
The proposed solution of using convolutional autoencoders for image compression and its resultant encryption demonstrates significant results in maintaining image quality while ensuring secure transfer. The paper demonstrates how convolutional autoencoders effectively reduce image size by encoding to latent space, thereby preventing unauthorised access during transmission. The architecture was designed with considerations for PSNR, SSIM, and MSE, which are metrics that gauge the quality of reconstructed images. The idea of using a specific batch size for final training yielded optimal results for stability and reconstruction quality. Additionally, employing a progressive approach with residual images allowed the model to better reconstruct the image with lower training times.
The scope for future work lies in the areas of quantization and entropy coding. Quantization involves mapping a range of values into a single value, while entropy coding relies on the occurrence of data and the number of bits required to represent them. Using the PyAC library for entropy coding were techniques used in few other papers and studies. Addressing these challenges and integrating these techniques could further enhance the efficiency and performance of the proposed autoencoder architecture.